Oriole Company bottles and distributes B-Lite, a diet soft drink. The beverage is sold for 60 cents per 16-ounce bottle to reYour answer is correct. Calculate variable cost per bottle. (Round variable cost per bottle to 3 decimal places, e.g. 0.251.)Compute the contribution margin ratio and the margin of safety ratio. (Round variable cost per bottle to 3 decimal places, e.
Answer #1

SOLUTION

Fixed expenses = Manufacturing overhead + Selling expenses + Administrative expenses

= 240,000 + 57,000 + 162,000 = 459,000

Variable expenses = Direct material + Direct labor + Manufacturing overhead + Selling expenses + Administrative expenses

= 500,000 + 390,000 + 410,000 + 107,000 + 21,000

= 1,428,000

Variable expense per unit = 1,428,000 / (2,040,000/60%)

= 1,428,000 / 3,400,000 = 0.42

Contribution margin per unit = 0.60 - 0.42 = 0.18

1. Break even point in units = Fixed expenses / Contribution margin per unit

= 459,000 / 0.18

= 2,550,000 units

Contribution margin ratio = 0.18 / 0.60 = 30%

Break even point in sales dollar = Fixed expenses / Contribution margin ratio

= 459,000 / 30% = 1,530,000

2. Contribution margin ratio = 0.18 / 0.60 = 30%

Margin of safety ratio = (Actual sales - Break even sales) / Actual sales * 100

= (2,040,000 - 1,530,000) / 2,040,000 * 100

= 25%

3. Required sales dollar = (Fixed expenses + Targeted income) / Contribution margin ratio

= (459,000 + 240,000) / 30%

= 699,000 / 30% = 2,330,000